# The $S_{n}$-equivariant Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$

### v1

@@ 2.3. $S_{2}\times S_{n}$-symmetric functions
[…]
@@
Applying Peter–Weyl theorem to $S_{2}\times S_{n}$-representations in $\mathsf{M}$, there exists a unique mixed Hodge structure $M_{\lambda,\nu}$ for each pair of partitions $\lambda\vdash 2$, $\nu\vdash n$, such that
@@
From [GK98, Proposition 8.10], for an $S_{2}$-representation $W$, we have
@@
where $D(-)$ is a linear operator defined on a basis by
@@
[…]
Symmetric functions associated to representations of the hyperoctahedral group $B_{n}:=S_{2}\wr S_{n}$ are needed for our study of decorated necklaces which arise in the localisation formula for ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. Our presentation is synthesised from that of Petersen [Pet12, §3] and Macdonald [Mac95, Appendix B].
@@
[…]
@@
For each ordered pair $(\lambda,\nu)$ of partitions such that $|\lambda|+|\nu|=n$, there is a unique irreducible representation $V_{\lambda,\nu}$ of $B_{n}$, and these form a complete list of the irreducible representations of $B_{n}$. For these representations, we have
@@
where $s_{\lambda},s_{\nu}\in\Lambda$ are the ordinary Schur functions and the maps $\iota_{\mathfrak{p}},\iota_{\mathfrak{q}}$ are as in (2.13). Now for $V$ is a $B_{n}$-representation in $\mathsf{M}$, the Peter–Weyl theorem for $\mathsf{M}$ allows us to define the enriched Frobenius characteristic of the mixed Hodge structure $V$ as
@@
[Pet12, Proposition 3.8] Let $H<B_{n}$ be a subgroup. Then
@@ 2.5. $S_{2}\times\mathbb{S}$ and $\mathbb{B}$-modules
In the following, we identify $B_{1}\cong S_{2}$ Applying the definition of plethysm given above, we find that
@@
$\mathfrak{p}_{1}\otimes_{S_{2}}\operatorname{ch}_{2,n}(V)=\operatorname{ch}_{n}(V_{1})+\operatorname{ch}_{n}(V_{2}),\quad\mathfrak{q}_{1}\circ_{S_{2}}\operatorname{ch}_{2,n}(V)=\operatorname{ch}_{n}(V_{1})-\operatorname{ch}_{n}(V_{2}).$
@@ 3. Composition and localisation on graded $\mathbb{S}$-spaces
In this section, we explain how the plethysm operations described in the previous section can be upgraded to compositions of (graded) $\mathbb{S}$-spaces. In the context of moduli spaces of marked nodal curves, the composition operation has the appealing geometric interpretation as attaching along marked points. This observation allows us to translate the geometric descriptions of strata in ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ into a plethystic formula for $\mathsf{b}_{1,r}$. This is the content of Proposition 3.9.
The composition operation on graded $\mathbb{S}$-spaces has been explored in [GP06], and we combine their formalism with torus localisation. In Lemma 3.2, we prove the (well-known) fact that the Frobenius characteristic of a graded $\mathbb{S}$-space with $\mathbb{C}^{\star}$-action can be recovered from the fixed locus. The only requirement is for the $\mathbb{C}^{\star}$-action to commute with the $S_{n}$-action; no smoothness or compactness property is needed.
@@ 3.1. $\mathbb{S}$-, $\mathbb{B}$-, and $S_{2}\times\mathbb{S}$-spaces
An $\mathbb{S}$-space $\mathcal{X}$ is a sequence $\mathcal{X}(n)$ of $S_{n}$-spaces, for each $n\geq 0$. $\mathbb{B}$-spaces and $S_{2}\times\mathbb{S}$ spaces are defined analogously, replacing $S_{n}$ with $B_{n}$ and $S_{2}\times S_{n}$, respectively. For each $\mathbb{A}\in\{\mathbb{S},\mathbb{B},S_{2}\times\mathbb{S}\}$, the Grothendieck groups $K_{0}(\mathsf{V},\mathbb{A})$, of $\mathbb{A}$-space can be made into a ring using the $\boxtimes$ product:
@@
where $G_{k}=S_{k}$ for all $k$ if $\mathbb{A}\in\{\mathbb{S},S_{2}\times\mathbb{S}\}$ or $G_{k}=B_{k}$ for all $k$ if $\mathbb{A}=\mathbb{B}$. In the case of $\mathbb{S}$-spaces $\mathcal{X}$ and $\mathcal{Y}$ which are moduli spaces of marked objects (e.g. pointed curves) the operation $\operatorname{Ind}_{S_{j}\times S_{n-j}}^{S_{n}}$ should be seen as taking all ways of distributing the $n$ marked points to the $j$ markings coming from $\mathcal{X}(j)$ and the remaining $n-j$ ones from $\mathcal{Y}(n-j)$.
@@
where $B_{k}$ acts diagonally. For $[\mathcal{X}]\in K_{0}(\mathsf{V},\mathbb{A})$ and $\mathbb{A}\in\{\mathbb{S},\mathbb{B},S_{2}\times\mathbb{S}\}$, we write $\mathsf{e}(\mathcal{X})=\sum_{n\geq 0}\mathsf{e}^{G_{n}}(\mathcal{X}(n))$ for the Serre characteristic of the $\mathbb{A}$-space $\mathcal{X}$; again $G_{k}$ is the appropriate group among $S_{n}$, $B_{n}$, or $S_{2}\times S_{n}$, and $\mathsf{e}$ takes values in $K_{0}(\mathsf{M})\otimes R$ where $R=\Lambda$, $R=\Lambda(S_{2})$, or $R=\Lambda_{2}\otimes\Lambda$, respectively. In each case, $\mathsf{e}$ is a ring homomorphism, which respects the composition structures.
@@ 3.2. Graded $\mathbb{S}$-spaces and $\mathbb{B}$-spaces
$[\mathcal{X}]\mapsto\sum_{d\geq 0}q^{d}\sum_{n\geq 0}[\mathcal{X}(n,d)]\in K_{0}(\mathsf{V},\mathbb{A})[[q]].$
@@
[tD87, §4.14] Let $G$ be a finite group and $H\subset G$ a subgroup. For an $H$-space $X$, there is an isomorphism $X/H\cong(\operatorname{Ind}_{H}^{G}X)/G$.
@@ 3.3. Localisation for graded $\mathbb{S}$-spaces
Working with each bidegree $(n,d)$, it suffices to show that if $X$ is a variety with commuting actions of $\mathbb{C}^{\star}$ and $S_{n}$, then $\chi^{S_{n}}(X)=\chi^{S_{n}}(X^{\mathbb{C}^{\star}})$. For this, suppose that $X\neq X^{\mathbb{C}^{\star}}$. We claim that it is possible to find an open subset $U$ of $X$ such that $U$ is preserved by both the $S_{n}$-action and by the $\mathbb{C}^{\star}$ action. We do this by first finding an open subset $U$ which is $\mathbb{C}^{\star}$-invariant such that the quotient $U\to U/\mathbb{C}^{\star}$ exists [Sum74] and then replacing $U$ by
$\bigcap_{\sigma\in S_{n}}\sigma(U),$
which is an $S_{n}$-invariant open subset. Possibly by again passing to an $S_{n}$-invariant open subset of the base, we can ensure that the $\mathbb{C}^{\star}$-action on $U$ is almost-free. Then by [GP06, Theorem 5.4], it follows that
$\chi^{S_{n}}(U)=\chi(\mathbb{C}^{\star})\cdot\chi^{S_{n}}(U/\mathbb{C}^{\star})=0.$
If $(X-U)^{\mathbb{C}^{\star}}=(X-U)$ then $X^{\mathbb{C}^{\star}}=X-U$, and $\chi^{S_{n}}(X)=\chi^{S_{n}}(X^{\mathbb{C}^{\star}})$. Otherwise, we can repeat this procedure by replacing $X$ by $X-U$. The proof is complete by Noetherian induction. ∎
@@ 3.4. Separating rational tails
We now show how to separate rational tails in the calculation of the class of the graded $\mathbb{S}$-space ¯$\overline{\mathcal{M}}_{g}(r)$. The first step in our calculation is to separate stable maps into:
@@
[…]
The core of $\mathbf{G}$, denoted as $\mathbf{G}^{\mathrm{core}}$ is the minimal $g$ subgraph of $\mathbf{G}$, together with a degree decoration restricted from $\mathbf{G}$, as well as one extra leg for each edge that connects $\mathbf{G}^{\mathrm{core}}$ to $\mathbf{G}\smallsetminus\mathbf{G}^{\mathrm{core}}$. A rational tail of $\mathbf{G}$ is a connected component of $\mathbf{G}\smallsetminus\mathbf{G}^{\mathrm{core}}$, where $\mathbf{G}$ and $\mathbf{G}^{\mathrm{core}}$ are topologized as 111-dimensional CW-complexes, ignoring legs and vertex decorations. Each rational tail carries degree decorations restricted from $\mathbf{G}$, and we endow each one with one extra leg for the edge in $\mathbf{G}$ that connects it to $\mathbf{G}^{\mathrm{core}}$.
@@
The following figure is a stable map dual graph $\mathbf{G}$ in ¯$\overline{\mathcal{M}}_{1,5}(\mathbb{P}^{r},7)$ and its core $\mathbf{G}^{\mathrm{core}}$ as well as rational tails $\mathbf{T}_{1},\dots,\mathbf{T}_{4}$. Note that in labeling the rational tails, we have made a non-canonical choice of ordering the tails.
@@
We denote by ¯$\overline{\mathcal{M}}_{g,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)\subset\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ the locally closed subscheme consisting of stable maps where the dual graph does not contain any rational tails. Equivalently, the dual graph of the stable map is equal to its own core.
[…]
@@
for $i=1,\dots,k$. The degrees $d^{\prime},d_{i}$ and valences $n^{\prime},n_{i}$ in general differ from $d$ and $n$ and can be determined from $\mathbf{G}$. Let $\star_{i}$ denote the extra leg in $\mathbf{T}_{i}$, and let $*_{i}$ denote the extra legs in $\mathbf{G}^{\mathrm{core}}$ connecting to $\mathbf{T}_{i}$. These legs define evaluation maps
@@
Define $\mathrm{Aut}(\hat{\mathbf{G}})$ to be the group of rooted tree automorphisms of $\hat{\mathbf{G}}$ that respect the decorations. It follows from the definitions that $\mathrm{Aut}(\hat{\mathbf{G}})\subset S_{k}$ is the Young subgroup that permutes blocks of subsets $\{\mathbf{T}_{i_{1}},\dots,\mathbf{T}_{i_{j}}\}$ where $\mathbf{T}_{i_{1}},\dots,\mathbf{T}_{i_{j}}$ are isomorphic, unmarked decorated trees. $\mathrm{Aut}(\hat{\mathbf{G}})$ thus naturally acts on $\mathcal{M}_{\mathbf{G}^{\mathrm{core}}}$ by permuting relevant markings and acts on $\prod_{i=1}^{k}\mathcal{M}_{\mathbf{T}_{i}}$ by permuting the entries. After describing the group action, we state the description of the stratum $\mathcal{M}_{\mathbf{G}}$ as follows.
@@
The stratum $\mathcal{M}_{\mathbf{G}}$ is isomorphic to the quotient ~$\widetilde{\mathcal{M}}_{\mathbf{G}}/\mathrm{Aut}(\hat{\mathbf{G}})$, where ~$\widetilde{\mathcal{M}}_{\mathbf{G}}$ is the fibre product
@@
There is an alternative description of ~$\widetilde{\mathcal{M}}_{\mathbf{G}}$ that will be useful. We note that $\mathrm{PGL}_{r+1}$ acts transitively on the target $\mathbb{P}^{r}$ and hence has an induced action on spaces of maps $\mathcal{M}_{\mathbf{T}_{i}}$, so that the evaluation maps $\mathrm{ev}_{\star_{i}}:\mathcal{M}_{\mathbf{T}_{i}}\to\mathbb{P}^{r}$ are $\mathrm{PGL}_{r+1}$-equivariant. The $\mathrm{PGL}_{r+1}$-equivariance implies that $\mathrm{ev}_{\star_{i}}$ are Zariski locally trivial fibrations. We define $\mathcal{M}^{\star}_{\mathbf{T}_{i}}:=\mathrm{ev}_{\star_{i}}^{-1}(p)$ to be the fibre over any $p\in\mathbb{P}^{r}$.
@@
The map ~$\widetilde{\mathcal{M}}_{\mathbf{G}}\to\mathcal{M}_{\mathbf{G}^{\mathrm{core}}}$ is Zariski locally trivial with fibres isomorphic to $\prod_{i=1}^{k}\mathcal{M}^{\star}_{\mathbf{T}_{i}}$.
@@
In the previous diagram, the vertical map $\prod_{i}\mathrm{ev}_{\star i}$ on the right is a Zariski locally trivial fibration with fibres $\prod_{i=1}^{k}\mathcal{M}^{\star}_{\mathbf{T}_{i}}$. Pulling back some Zariski local trivialisation of $\prod_{i}\mathrm{ev}_{\star i}$ along the fibre diagram trivialises the vertical map ~$\widetilde{\mathcal{M}}_{\mathbf{G}}\to\mathcal{M}_{\mathbf{G}^{\mathrm{core}}}$ on the left with the same fibres. ∎
@@
Let ¯$\overline{\mathcal{M}}_{g}(r)$ be the graded $\mathbb{S}$-space given by $(n,d)\mapsto\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$, and let ¯$\overline{\mathcal{M}}^{\mathrm{nrt}}_{g}(r)$ be the graded $\mathbb{S}$-space given by $(n,d)\mapsto\overline{\mathcal{M}}_{g,n}^{{\mathrm{nrt}}}(\mathbb{P}^{r},d)$. As earlier, the Frobenius characteristics of these spaces are denoted as $\mathsf{b}_{g,r}$ and $\mathsf{b}_{g,r}^{\mathrm{nrt}}$, respectively.
@@
be a fibre of the map, which is an $S_{n}$-space. Define ¯$\overline{\mathcal{M}}^{\star}_{0}(r)$ as the graded $\mathbb{S}$-space given by
@@
[¯$[\overline{\mathcal{M}}^{\star}_{0}(r)]=\frac{D\overline{\mathcal{M}}_{0}(r)}{[\mathbb{P}^{r}]},$
@@
The following is the main result of the section. It states how the Serre characteristic of ¯$\overline{\mathcal{M}}_{1}(r)$ is determined by those of the no-rational-tails locus ¯$\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)$ and the space ¯$\overline{\mathcal{M}}^{\star}_{0}(r)$ of maps from rational tails.
@@
¯$\overline{\mathcal{M}}_{1}(r)=\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)\circ\left([\operatorname{Spec}\mathbb{C}]+\frac{D\overline{\mathcal{M}}_{0}(r)}{[\mathbb{P}^{r}]}\right).$

### v2

@@ 2.3. $S_{2}\times S_{n}$-symmetric functions
Consider the tensor product $K_{0}(\mathsf{M})\otimes\Lambda\otimes\Lambda$, thought of as a ring of bisymmetric functions. The characteristics associated to $S_{2}\times\mathbb{S}$-spaces generate the vector subspace $K_{0}(\mathsf{M})\otimes\Lambda_{2}\otimes\Lambda$ spanned by bihomogeneous polynomials of bidegrees $(2,i)$ for all $i\geq 0$. There are two convenient linear bases for $\Lambda_{2}$ that we will use: the first is
@@
Applying the Peter–Weyl theorem to $S_{2}\times S_{n}$-representations in $\mathsf{M}$, there exists a unique mixed Hodge structure $M_{\lambda,\nu}$ for each pair of partitions $\lambda\vdash 2$, $\nu\vdash n$, such that
@@
From [21, Proposition 8.10], for an $S_{2}$-representation $W$, we have
@@
where D()D(-) is a linear operator defined on a basis by
@@
2.4. Type-$B$ symmetric functions
Symmetric functions associated to representations of the hyperoctahedral group $B_{n}:=S_{2}\wr S_{n}$ are needed for our study of decorated necklaces which arise in the localisation formula for $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. Our presentation is synthesised from that of Petersen [54, §3] and Macdonald [44, 43, Appendix B].
@@ 2.4. Type-$B$ symmetric functions
Given a $B_{n}$-representation $V$ in $\mathsf{Vect}_{\mathbb{Q}}$, the type-$B$ Frobenius characteristic is given by
@@
For each ordered pair $(\lambda,\nu)$ of partitions such that $|\lambda|+|\nu|=n$, there is a unique irreducible representation $V_{\lambda,\nu}$ of $B_{n}$, and these form a complete list of the irreducible representations of $B_{n}$ [43, §1.B.9]. For these representations, we have
@@
where $s_{\lambda},s_{\nu}\in\Lambda$ are the ordinary Schur functions and the maps $\iota_{\mathfrak{p}},\iota_{\mathfrak{q}}$ are as in (2.13). Now for $V$ a $B_{n}$-representation in $\mathsf{M}$, the Peter–Weyl theorem for $\mathsf{M}$ allows us to define the enriched Frobenius characteristic of the mixed Hodge structure $V$ as
@@
[54, Proposition 3.8] Let $H<B_{n}$ be a subgroup. Then
@@ 2.5. $S_{2}\times\mathbb{S}$ and $\mathbb{B}$-modules
In the following, we identify $B_{1}\cong S_{2}$ . Applying the definition of plethysm given above, we find that
@@
$\mathfrak{p}_{1}\circ_{S_{2}}\operatorname{ch}_{2,n}(V)=\operatorname{ch}_{n}(V_{1})+\operatorname{ch}_{n}(V_{2}),\quad\mathfrak{q}_{1}\circ_{S_{2}}\operatorname{ch}_{2,n}(V)=\operatorname{ch}_{n}(V_{1})-\operatorname{ch}_{n}(V_{2}).$
@@ 3. Composition and localisation on graded $\mathbb{S}$-spaces
In this section, we explain how the plethysm operations described in the previous section can be upgraded to compositions of (graded) $\mathbb{S}$-spaces. In the context of moduli spaces of marked nodal curves, the composition operation has the appealing geometric interpretation as attaching along marked points. This observation allows us to translate the geometric descriptions of strata in $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ into a plethystic formula for $\mathsf{b}_{1,r}$. This is the content of Proposition 3.9.
The composition operation on graded $\mathbb{S}$-spaces has been explored in [25], and we combine their formalism with torus localisation. In Lemma 3.2, we prove the (well-known) fact that the Frobenius characteristic of a graded $\mathbb{S}$-space with $\mathbb{C}^{\star}$-action can be recovered from the fixed locus. The only requirement is for the $\mathbb{C}^{\star}$-action to commute with the $S_{n}$-action; no smoothness or compactness property is needed.
@@ 3.1. $\mathbb{S}$-, $\mathbb{B}$-, and $S_{2}\times\mathbb{S}$-spaces
An $\mathbb{S}$-space $\mathcal{X}$ is a sequence $\mathcal{X}(n)$ of $S_{n}$-spaces, for each $n\geq 0$. $\mathbb{B}$-spaces and $S_{2}\times\mathbb{S}$ spaces are defined analogously, replacing $S_{n}$ with $B_{n}$ and $S_{2}\times S_{n}$, respectively. For each $\mathbb{A}\in\{\mathbb{S},\mathbb{B},S_{2}\times\mathbb{S}\}$, the Grothendieck groups $K_{0}(\mathsf{V},\mathbb{A})$ of $\mathbb{A}$-spaces can be made into a ring using the $\boxtimes$ product:
@@
where $G_{k}=S_{k}$ for all $k$ if $\mathbb{A}\in\{\mathbb{S},S_{2}\times\mathbb{S}\}$ or $G_{k}=B_{k}$ for all $k$ if $\mathbb{A}=\mathbb{B}$. In the case of $\mathbb{S}$-spaces $\mathcal{X}$ and $\mathcal{Y}$ which are moduli spaces of marked objects (e.g. pointed curves) the operation $\operatorname{Ind}_{S_{j}\times S_{n-j}}^{S_{n}}$ should be seen as taking all ways of distributing the $n$ marked points to the $j$ markings coming from $\mathcal{X}(j)$ and the remaining njn-j ones from $\mathcal{Y}(n-j)$.
@@
where $B_{k}$ acts diagonally. For $\mathcal{X}]\in K_{0}(\mathsf{V},\mathbb{A})$ and $\mathbb{A}\in\{\mathbb{S},\mathbb{B},S_{2}\times\mathbb{S}\}$, we write
$\mathsf{e}(\mathcal{X})=\sum_{n\geq 0}\mathsf{e}^{G_{n}}(\mathcal{X}(n))$
for the Serre characteristic of the $\mathbb{A}$-space $\mathcal{X}$; again $G_{k}$ is the appropriate group among $S_{n}$, $B_{n}$, or $S_{2}\times S_{n}$, and $\mathsf{e}$ takes values in $K_{0}(\mathsf{M})\otimes R$ where $R=\Lambda$, $R=\Lambda(S_{2})$, or $R=\Lambda_{2}\otimes\Lambda$, respectively. In each case, $\mathsf{e}$ is a ring homomorphism, which respects the composition structures.
@@ 3.2. Graded $\mathbb{S}$-spaces and $\mathbb{B}$-spaces
$\mathcal{X}]\mapsto\sum_{d\geq 0}q^{d}\sum_{n\geq 0}[\mathcal{X}(n,d)]\in K_{0}(\mathsf{V},\mathbb{A})[[q]].$
@@
[64, §4.14] Let $G$ be a finite group and $H\subset G$ a subgroup. For an $H$-space $X$, there is an isomorphism $X/H\cong(\operatorname{Ind}_{H}^{G}X)/G$.
@@ 3.3. Localisation for graded $\mathbb{S}$-spaces
Working with each bidegree $(n,d)$, it suffices to show that if $X$ is a variety with commuting actions of $\mathbb{C}^{\star}$ and $S_{n}$, then $\chi^{S_{n}}(X)=\chi^{S_{n}}(X^{\mathbb{C}^{\star}})$. By the Lefschetz fixed-point theorem, we have
$\chi^{S_{n}}(X)=\frac{1}{n!}\sum_{\sigma\in S_{n}}\chi(X^{\sigma})\prod_{i=1}^{n}p_{i}^{\lambda_{i}(\sigma)}$
where $\lambda_{i}(\sigma)$ is the number of $i$-cycles in $\sigma$. Since the $S_{n}$-action commutes with the $\mathbb{C}^{\star}$-action, we have
$(X^{\sigma})^{\mathbb{C}^{\star}}=(X^{\mathbb{C}^{\star}})^{\sigma}$
for all $\sigma\in S_{n}$. Therefore, to prove the lemma, it suffices to prove that for all $\sigma\in S_{n},$
$\chi(X^{\sigma})=\chi\left((X^{\sigma})^{\mathbb{C}^{\star}}\right).$
To show this, we will prove that for an arbitrary $\mathbb{C}^{\star}$-variety $Y$, we have $\chi(Y)=\chi(Y^{\mathbb{C}^{\star}})$, which is likely well-known to experts. Let $Y^{sing}\subset Y$ denote the singular locus of $Y$, and let $Y^{sm}=Y\smallsetminus Y^{sing}$ be the smooth locus of $Y^{sm},$ which is open in $Y.$ Since $\mathbb{C}^{\star}$ must preserve $Y^{sing}$, both $Y^{sing}$ and $Y^{sm}$ are $\mathbb{C}^{\star}$-invariant, and we have a partition
$Y^{\mathbb{C}^{\star}}=(Y^{sm})^{\mathbb{C}^{\star}}\sqcup(Y^{sing})^{\mathbb{C}^{\star}}.$
By induction on dimension, we can assume that $\chi((Y^{sing})^{\mathbb{C}^{\star}})=\chi(Y^{sing})$, so we are reduced to showing that $\chi(Y^{sm})=\chi((Y^{sm})^{\mathbb{C}^{\star}})$. Since $\mathbb{C}^{\star}$ is connected, each $\mathbb{C}^{\star}$-orbit is irreducible and is contained in some irreducible component of $Y^{sm}.$ Therefore, the $\mathbb{C}^{\star}$-action preserves each irreducible component of $Y^{sm},$ so we reduce to the case where $Y^{sm}$ is irreducible.
In this case, we can find a $\mathbb{C}^{\star}$-invariant open subset $U\subset Y$ [63]. By passing further to an open subset if necessary, we may assume that $\mathbb{C}^{\star}$-action on $U$ is almost free, so that the quotient $U\to U/\mathbb{C}^{\star}$ exists, and in particular is étale locally a $\mathbb{C}^{\star}$-bundle. It follows that $\chi(U)=0$. If $(Y^{sm}\smallsetminus U)^{\mathbb{C}^{\star}}=Y^{sm}\smallsetminus U$, then $(Y^{sm})^{\mathbb{C}^{\star}}=Y^{sm}\smallsetminus U$ and $\chi(Y^{sm})=\chi((Y^{sm})^{\mathbb{C}^{\star}})$. Otherwise, we can repeat this procedure by replacing $Y^{sm}$ with $Y^{sm}\smallsetminus U$. The proof is complete by Noetherian induction. ∎
@@ 3.4. Separating rational tails
We now show how to separate rational tails in the calculation of the class of the graded $\mathbb{S}$-space $\overline{\mathcal{M}}_{g}(r)$. The first step in our calculation is to separate stable maps into:
@@
Let $f:C\to\mathbb{P}^{r}]\in\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ be a genus-$g$, $n$-pointed stable map. Let $\mathbf{G}$ be the dual graph of $f$; this is the dual graph of the curve $C$, decorated additionally by the degree of $f$ along each component. We refer to such graphs as $(g,n,d)$-graphs in [37, Definition 1.1].
The core of $\mathbf{G}$, denoted as $\mathbf{G}^{\mathrm{core}}$ is the minimal genus-$g$ subgraph of $\mathbf{G}$, together with a degree decoration restricted from $\mathbf{G}$, as well as one extra leg for each edge that connects $\mathbf{G}^{\mathrm{core}}$ to $\mathbf{G}\smallsetminus\mathbf{G}^{\mathrm{core}}$. A rational tail of $\mathbf{G}$ is a connected component of $\mathbf{G}\smallsetminus\mathbf{G}^{\mathrm{core}}$, where $\mathbf{G}$ and $\mathbf{G}^{\mathrm{core}}$ are topologized as $1$-dimensional CW-complexes, ignoring legs and vertex decorations. Each rational tail carries degree decorations restricted from $\mathbf{G}$, and we endow each one with one extra leg for the edge in $\mathbf{G}$ that connects it to $\mathbf{G}^{\mathrm{core}}$.
@@
The following figure is a stable map dual graph $\mathbf{G}$ in $\overline{\mathcal{M}}_{1,5}(\mathbb{P}^{r},7)$ and its core $\mathbf{G}^{\mathrm{core}}$ as well as rational tails $\mathbf{T}_{1},\dots,\mathbf{T}_{4}$. Note that in labeling the rational tails, we have made a non-canonical choice of ordering the tails.
@@
We denote by $\overline{\mathcal{M}}_{g,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)\subset\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ the locally closed subscheme consisting of stable maps where the dual graph does not contain any rational tails. Equivalently, the dual graph of the stable map is equal to its own core.
Before stating the main formula of the section, we see how the above definitions allow us to express each stratum in $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ as a fibre product. Let $\mathbf{G}$ be an $(g,n,d)$-graph, so that it specifies a locally closed stratum $\mathcal{M}_{\mathbf{G}}\subset\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$. Let $\mathbf{G}^{\mathrm{core}}$ be the core, and let $\mathbf{T}_{1},\dots,\mathbf{T}_{k}$ be the rational tails of $\mathbf{G}$. The graphs $\mathbf{G}^{\mathrm{core}}$ and $\mathbf{T}_{1},\dots,\mathbf{T}_{k}$ specify strata
@@
for $i=1,\dots,k$. The degrees $d^{\prime},d_{i}$ and valences $n^{\prime},n_{i}$ in general differ from $d$ and $n$ and can be determined from $\mathbf{G}$. Let $\star_{i}$ denote the extra leg in $\mathbf{T}_{i}$, and let $_{i}$ denote the extra legs in $\mathbf{G}^{\mathrm{core}}$ connecting to $\mathbf{T}_{i}$. These legs define evaluation maps
@@
[…]
@@
$\widetilde{\mathcal{M}}_{\mathbf{G}}$$\prod_{i=1}^{k}\mathcal{M}_{\mathbf{T}_{i}}$$\mathcal{M}_{\mathbf{G}^{\mathrm{core}}}$$\prod_{i=1}^{k}\mathbb{P}^{r}$$\square$$\scriptstyle{\prod_{i}\mathrm{ev}_{\star_{i}}}$$\scriptstyle{\prod_{i}\mathrm{ev}_{*_{i}}}$
There is an alternative description of $\widetilde{\mathcal{M}}_{\mathbf{G}}$ that will be useful. We note that $\mathrm{PGL}_{r+1}$ acts transitively on the target $\mathbb{P}^{r}$ and hence has an induced action on spaces of maps $\mathcal{M}_{\mathbf{T}_{i}}$, so that the evaluation maps $\mathrm{ev}_{\star_{i}}:\mathcal{M}_{\mathbf{T}_{i}}\to\mathbb{P}^{r}$ are $\mathrm{PGL}_{r+1}$-equivariant. The $\mathrm{PGL}_{r+1}$-equivariance implies that $\mathrm{ev}_{\star_{i}}$ are Zariski locally trivial fibrations. We define $\mathcal{M}^{\star}_{\mathbf{T}_{i}}:=\mathrm{ev}_{\star_{i}}^{-1}(p)$ to be the fibre over any $p\in\mathbb{P}^{r}$.
@@
The map $\widetilde{\mathcal{M}}_{\mathbf{G}}\to\mathcal{M}_{\mathbf{G}^{\mathrm{core}}}$ is Zariski locally trivial with fibres isomorphic to $\prod_{i=1}^{k}\mathcal{M}^{\star}_{\mathbf{T}_{i}}$.
@@
In the previous diagram, the vertical map $\prod_{i}\mathrm{ev}_{\star i}$ on the right is a Zariski locally trivial fibration with fibres $\prod_{i=1}^{k}\mathcal{M}^{\star}_{\mathbf{T}_{i}}$. Pulling back some Zariski local trivialisation of $\prod_{i}\mathrm{ev}_{\star i}$ along the fibre diagram trivialises the vertical map $\widetilde{\mathcal{M}}_{\mathbf{G}}\to\mathcal{M}_{\mathbf{G}^{\mathrm{core}}}$ on the left with the same fibres. ∎
@@
Let $\overline{\mathcal{M}}_{g}(r)$ be the graded $\mathbb{S}$-space given by $(n,d)\mapsto\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$, and let $\overline{\mathcal{M}}^{\mathrm{nrt}}_{g}(r)$ be the graded $\mathbb{S}$-space given by $(n,d)\mapsto\overline{\mathcal{M}}_{g,n}^{{\mathrm{nrt}}}(\mathbb{P}^{r},d)$. As earlier, the Frobenius characteristics of these spaces are denoted as $\mathsf{b}_{g,r}$ and $\mathsf{b}_{g,r}^{\mathrm{nrt}}$, respectively.
@@
be a fibre of the map, which is an $S_{n}$-space. Define $\overline{\mathcal{M}}^{\star}_{0}(r)$ as the graded $\mathbb{S}$-space given by
@@
$\overline{\mathcal{M}}^{\star}_{0}(r)]=\frac{D\overline{\mathcal{M}}_{0}(r)}{[\mathbb{P}^{r}]},$
@@
The following is the main result of the section. It states how the Serre characteristic of $\overline{\mathcal{M}}_{1}(r)$ is determined by those of the no-rational-tails locus $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)$ and the space $\overline{\mathcal{M}}^{\star}_{0}(r)$ of maps from rational tails.
@@
$\overline{\mathcal{M}}_{1}(r)=\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)\circ\left([\operatorname{Spec}\mathbb{C}]+\frac{D\overline{\mathcal{M}}_{0}(r)}{[\mathbb{P}^{r}]}\right).$